OpenStudy (swissgirl):

Solve for x: 3^x=x^2

OpenStudy (cwrw238):

the only way i can think of to solve this is to draw graphs of 3^x and x^2 and find where they intersect

OpenStudy (apoorvk):

\(3^x = x^2\) \(=> \log_3(x^2)= x\) \(=> 2\log_3x = x\) Now what?? Aww woof woof!!!! Wolf says: http://www.wolframalpha.com/input/?i=3%5Ex+%3D+x%5E2

OpenStudy (apoorvk):

or the graph. -_-